OpenStudy (anonymous):

Find the value of 'k' that satisfies the following equation: -5k+4+8k-1=-3k+3-2k+1

OpenStudy (whpalmer4):

\[-5k+4+8k-1=-3k+3-2k+1\] What happens if you collect all of the like terms together?

OpenStudy (anonymous):

-5k+4+8k-1=-3k+3-2k+1 i.e. 3k+3= -5k +4 i.e. 8K = 4-3 i.e. 8k= 1 i.e. K= 1/8

OpenStudy (anonymous):

Wow thanks to both ! Now i can understand to solve problems like this